OpenStudy (mmend98):

The average kinetic energy of 1 mole of a gas at -32 degrees Celsius is: 3.01 x 103 J 3.99 x 102 J -3.99 x 102 J 3.80 x 103 J

OpenStudy (taramgrant0543664):

KE=(3/2)nRT Where KE is kinetic energy n is number of moles R is 8.31J/K T is temperature in Kelvin

OpenStudy (mmend98):

What is the final answer? its okay this isn't a test. ill give u a medal lol @taramgrant0543664

OpenStudy (taramgrant0543664):

KE=(3/2)(1.0)(8.31)(-32+273) KE=3004J KE=3.00x10^3J

OpenStudy (taramgrant0543664):

If you use 273.15 then KE=3005 so rounding that to three significant digits it would be 3010 making it 3.01x10^3 in scientific notation